Q1. THE FOLLOWING FUNCTIONS ARE GIVEN FOR AN ECONOMY:-
C = 100+ 0.9 YD
I = 600- 30i
T= 1/3 Y
G= 300
MD = 0.4 Y – 50 i
MS = 1040 (NOMINAL MONEY SUPPLY)
P = 2
1. Calculate equilibrium income and rate of interest.
2. Find out the value of simple multiplier.
3. If govt increases its expenditure by 100 units, then what would be the increase in GDP?
1
Expert's answer
2020-05-18T11:36:27-0400

1)

"MD=\frac{MS}{P}"

"0.4Y-50i=\frac{1040}{2}"

"i=\frac{0.4Y-520}{50}"

"Y=AD"

"Y=100+0.9(Y-\frac{1}{3}Y)+600+30\frac{0.4Y-520}{50}+300"

"Y=1000+0.9Y-0.3Y+\frac{1.2Y-1560}{5}"

"Y=1000+0.6Y+0.24Y-312"

"Y=688+0.84Y"

"0.16Y=688"

"Y=\frac{688}{0.16}"

"Y=4300"

"i=\frac{0.4*4300-520}{50}=24"

2)

"m=\frac{1}{0.16}=6.25"

3)

"Y=\frac{688+100}{0.16}=4925"